OpenStudy (anonymous):

Given the function f defined by f(x)=x^2-2x-3, find the equation of the tangent line to the curve at x=3.

OpenStudy (bahrom7893):

Find first derivative then plug in x = 3 to find the slope..

OpenStudy (bahrom7893):

f' = 2x - 2 f'(3) = 2*3-2 = 4

OpenStudy (bahrom7893):

So slope of tangent line is M = 4

OpenStudy (bahrom7893):

now u need x = 3 and f(3) = 3^2 - 2*3 - 3 = 9 - 6 - 3 = 0 So u have all u need: x = 3; y = 0; M = 4

OpenStudy (bahrom7893):

y - 0 = 4(x-3)

OpenStudy (bahrom7893):

y = 4x - 12

OpenStudy (bahrom7893):

Please click on become a fan if I helped, I really want to get to the next level!! Thanks =)

OpenStudy (anonymous):

Yes. I am understanding this now. These problems are from my old 1972 calculus book.